OpenStudy (anonymous):

Simplify sqrt72 - sqrt147- 5sqrt8 + 3sqrt48

OpenStudy (anonymous):

options are -2sqrt35 sqrt5 -8 + 5sqrt6 -4sqrt2 + 5sqrt3

OpenStudy (ghazi):

2\[2\sqrt{18}- \sqrt{147}-5\sqrt{8}+3*4\sqrt{3}\]

OpenStudy (ghazi):

\[\sqrt{147}= 7\sqrt{3}\]

OpenStudy (anonymous):

thanks but.... that's not one of my options

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

i see waht u did

OpenStudy (anonymous):

nvm im confused can you explain please

OpenStudy (ghazi):

i have just reduced now you can simplify

OpenStudy (ghazi):

\[2\sqrt{18}=2*3\sqrt{2}\]=\[6\sqrt{2}\]

OpenStudy (ghazi):

\[5\sqrt{8}=5*2\sqrt{2}=10\sqrt{2}\]

OpenStudy (ghazi):

finally you'll get \[6\sqrt{2}-7\sqrt{3}-10\sqrt{2}+4\sqrt{3}= -4\sqrt{2}-3\sqrt{3}=-(4\sqrt{2}+3\sqrt{3})\]

OpenStudy (ghazi):

hope you're done now

OpenStudy (anonymous):

i got -4sqrt2 + 5sqrt3

OpenStudy (ghazi):

nope it's -4sqrt2-3sqrt3

OpenStudy (anonymous):

no uts 5 sqrt 3 you were supposed to add 3 sqrt 3

OpenStudy (anonymous):

i mean 3 sqrt 48

OpenStudy (ghazi):

okay...